An airline has requests for standby flights at half of the usual oneway air fare. Past experience has shown that these passengers have about a 1 in 5 chance of getting on the standby flight. When they fail to get on a flight as a standby, the only other choice is to fly first class on the next flight out. Suppose that the usual one way air fare to a certain city is $100 and the cost of flying first class is $395. Should a passenger who wishes to fly to this city opt to fly as a standby? [Hint Find the expected cost of the trip for a person flying standby.]
Answer #1

Usual one way air fare =$100. So, Standby air fare =0.5(100) =$50

Expected cost =1/5($50) + 4/5($395) =$10 + $316 =$326

No. A passenger who wishes to fly to this city should not opt to fly as a standby because the expected cost of $326 is greater than the usual one way air fare of $100.